OpenStudy (anonymous):

"how" to integrate: 1/(sinx + cosx)^4

OpenStudy (anonymous):

accidentally closed this question :/

OpenStudy (anonymous):

@amistre64 you know?

OpenStudy (amistre64):

it does seem to be a bit peculiar; im wondering how a u sub would look

OpenStudy (amistre64):

u = sinx + cosx du = cosx - sinx dx that doesnt looks to useful to me

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+1%2F%28sinx%2Bcosx%29 that looks nightmarish

OpenStudy (amistre64):

forgot the ^4

OpenStudy (amistre64):

thats better

OpenStudy (anonymous):

tried that to a certain extend already, leads me no where (at least for me anyways...) also, wolframalpha has provided a u sub, where u=tan(y/2) or something like that, but the solution is extremely non-practical....and ofc I'm assuming a better method exist...

OpenStudy (amistre64):

well, i can think of power series, and euler forms

OpenStudy (anonymous):

i can do ^1 and ^2, but ^4 is a real pain

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+1%2F%28sinx%2Bcosx%29%5E4 take the derivative of that to see if you can see a clue as to how to reverse it for integration

OpenStudy (anonymous):

just when you are happily doing a double integral, you come to a single variable integral you can't do....smh

OpenStudy (amistre64):

can we convert it to a double? or a polar?

OpenStudy (anonymous):

already tried wolframaphla :(, non-practical solution

OpenStudy (amistre64):

can we say: u = sinx v = cosx and do a multivarable integartion?

OpenStudy (tkhunny):

You can ALWAYS use the the tan(y/2), but seriously? Why not simplify your life? \(\sin(x) + \cos(x) = \sqrt{2}\cdot\cos(x - \pi/4) \)

OpenStudy (amistre64):

... not that i can type it :/

OpenStudy (amistre64):

that was one form i never learnt :)

OpenStudy (anonymous):

lol, never seem that form as well, gona try it now....